# Predictability Awareness For Efficient and Robust Multi-Agent Coordination

### v1

@@ 1. Introduction
Receding Horizon Trajectory Optimization allows for flexible and anticipative planning while ensuring compliance with e.g. safety constraints in multi-agent navigation problems. However, planning a trajectory that explicitly accounts for interactions among agents generally requires solving a joint optimization problem. A variety of joint planning methods can be found in literature, e.g. Huang et al. (2024); Chen et al. (2023), of which game theoretic approaches best capturing agent interaction complexities Wang et al. (2022). By modelling other agents as rational actors, game theoretic approaches cast the joint optimization as a constrained dynamic game and seek to find equilibrium solutions. Although this often results in stable and coordinated interactions Fisac et al. (2018); Cleac’h et al. (2020), game theoretic approaches suffer from the curse of dimensionality, as the planning complexity grows exponentially with the number of agents Schwarting et al. (2018). Additionally, modelling other agents as rational is a strong assumption which will not hold in practice, especially when interacting with human agents Colman (2003); Bossaerts and Murawski (2017).
@@ 2. Related Work
In the field of Human-Robot Interaction, legibility and predictability of motion have been studied to improve coordination by designing agent behaviors that clearly communicate intention and avoid surprising observers Dragan et al. (2013). Often both objectives overlap Bhatt et al. (2023). Traditional formulations of this problem are not well suited for receding horizon applications as they optimize over complete trajectories and rely on utility-based analytical models of observer expectations Dragan and Srinivasa (2014). Additionally, the observer is modeled as inactive, thus having no influence on the planning agent. This assumption breaks down in multi-agent navigation where the observer and the agent share the workspace and influence each other. Several works have explored the adaptation of these concepts to an interactive multi-agent context. Bastarache et al. (2023) define dynamic goal regions around neighboring agents and optimize for reduced uncertainty about the collision avoidance strategy. Geldenbott and Leung (2024) show how increasing action penalties at later horizon steps causes agents to more rapidly demonstrate their avoidance strategy. This accelerates intent inference giving agents better anticipation. Brüdigam et al. (2018) defines hand-crafted legibility costs for planning in highway driving. These methods are often designed to target a specific type of observer model. In contrast, our approach minimizes a predictability surrogate that allows modeling the observers with an arbitrary prediction model choice.
@@ 4.1. Derivation of a predictability aware cost function
where $\mathbb{KL}$ denotes the KL-Divergence. The right-hand side provides an upper bound on the free energy, and one can minimise this instead of the free energy. It resembles a standard control objective, and the terms allow for good conceptual understanding of the effect they have: A Performance Cost and Predictability Cost respectively, which penalizes agents for acting unpredictably. Using this newly found expression as a stage cost, we can craft the following cost function as a stage cost for a planning problem:
@@
We can emphasize now the intuition behind using the free energy as a way of incorporating predictability into optimal control. Eq. (4) is minimised precisely when $\mathcal{Q}=\mathcal{Q}^{*}=\mathcal{P}$. That is, the trajectory distribution executed is exactly the optimal cost trajectory distribution, and this matches the predicted distribution. Under this condition, the agent is behaving without surprising external observers and simultaneously obtaining optimal cost in its objective.
@@ 4.2. Integration with a Planner and Practicalities
$J(\tau_{0:K})=\sum_{k=0}^{K}J_{k}(\boldsymbol{x}_{k},\boldsymbol{u}_{k})+\gamma^{t}\lambda\mathbb{KL}(q(\boldsymbol{x}_{k})||p(\boldsymbol{x}_{k})).$ (5)
@@ 5.1. Planner
For all experiments in this section, we use a sampling-based planner, namely Model Predictive Path Integral (MPPI) control, based on the methodology presented in Williams et al. (2017b). MPPI places no restrictions on dynamics model or cost function and converges well toward optima with a moderate amount of samples Williams et al. (2017a). Given a nominal control sequence as an initial guess, MPPI applies Gaussian noise at each step to generate a set of $M$ control sequence samples. It then uses a state transition function $f(\cdot)$ to simulate their corresponding $M$ state trajectories. Each of the resulting state trajectories is evaluated based on the cost defined in (5), resulting in a total sample cost $J_{m}$. Once $J_{m}$, $\forall m\in[1,...,M]$ is computed, importance sampling weights, $w_{m}$, can be calculated as:
@@ 5.4.2. Robot-Robot Traffic Scenario
When agents fail to coordinate in road scenarios, they often experience deadlocks or, in the worst case, collisions. In Figure 3(a), an example of a deadlock is illustrated. Deadlocks are common in limited space environments such as intersections or narrow passages. Initially, the model may predict one agent will yield while the other advances. However, as deviations occur and both agents hesitate, their predictions begin to reinforce each other’s hesitation, creating the deadlock. The model may then be unable to introduce asymmetry to prioritize one of the agents in ambiguous situations, preventing the agents from breaking away from the deadlock. Results show that agents incorporating predictability into their models achieve better coordination, as indicated by less pronounced slowdowns resulting in higher travelled distance and the disappearance of deadlocks as seen in Table 2. When examining other metrics, the benefits of incorporating predictability are not as pronounced, especially for higher $\lambda$. This occurs because the prediction model is not explicitly conditioned to align with the road geometry (Figures 3(c),3(d)). Since the planner is required to track a reference path, deviations between the predictions and the reference path can push the agent to deviate from the path, requiring small adjustments more frequently for higher higher $\lambda$. Although this problem has marginal impact on the overall performance of the agent, the reduction in planning effort may be mitigated.
@@ Discussion
Balancing predictability and performance cost, determined by $\lambda$, is complex and context-dependent. Dynamically adjusting $\lambda$ as agents interact could improve performance, increasingly prioritizing predictability in safety-critical moments. Developing adaptive heuristics for this adjustment, as suggested by previous work Bastarache et al. (2023); Dragan and Srinivasa (2014), would be a valuable research direction. The free energy framework in Section 4 could be extended to dynamically adjust prediction distributions over time in reaction to the samples, enhancing anticipation and balancing. However, continuous updates bring computational challenges, especially in multi-agent systems, making scalability difficult. Simplifications or assumptions may help overcome these challenges, merging predictability and legibility to boost performance.

### v2

@@ 1. Introduction
footnotetext: Equal contribution.
This research is supported by funding from the Dutch Research Council NWO-NWA, within the “Acting under Uncertainty” (ACT) project (Grant No.NWA.1292.19.298). Work done partially while at Delft University of Technology. Author acknowledges partial support from UKRI grant EP/W002949/1
Receding Horizon Trajectory Optimization allows for flexible and anticipative planning while ensuring compliance with e.g. safety constraints in multi-agent navigation problems. However, planning a trajectory that explicitly accounts for interactions among agents generally requires solving a joint optimization problem. A variety of joint planning methods can be found in literature, e.g. Huang et al. (2024); Chen et al. (2023), of which game theoretic approaches best capturing agent interaction complexities Wang et al. (2022). By modeling other agents as rational actors, game theoretic approaches cast the joint optimization as a constrained dynamic game and seek to find equilibrium solutions. Although this often results in stable and coordinated interactions Fisac et al. (2018); Cleac’h et al. (2020), game theoretic approaches suffer from the curse of dimensionality, as the planning complexity grows exponentially with the number of agents Schwarting et al. (2018). Additionally, modeling other agents as rational is a strong assumption which will not hold in practice, especially when interacting with human agents Colman (2003); Bossaerts and Murawski (2017).
@@ 2. Related Work
In the field of Human-Robot Interaction, legibility and predictability of motion have been studied to improve coordination by designing agent behaviors that clearly communicate intention and avoid surprising observers Dragan et al. (2013). Often both objectives overlap Bhatt et al. (2023). Traditional formulations of this problem are not well suited for receding horizon applications as they optimize over complete trajectories and rely on utility-based analytical models of observer expectations Dragan and Srinivasa (2014). Additionally, the observer is modeled as inactive, thus having no influence on the planning agent. This assumption breaks down in multi-agent navigation where the observer and the agent share the workspace and influence each other. Several works have explored the adaptation of these concepts to an interactive multi-agent context. In single agent RL settings, the impact of predictability objectives has been recently studied in Ornia et al. (2023). In multi-agent scenarios, Bastarache et al. (2023) define dynamic goal regions around neighboring agents and optimize for reduced uncertainty about the collision avoidance strategy. Geldenbott and Leung (2024) show how increasing action penalties at later horizon steps causes agents to more rapidly demonstrate their avoidance strategy. This accelerates intent inference giving agents better anticipation. Brüdigam et al. (2018) defines hand-crafted legibility costs for planning in highway driving. These methods are often designed to target a specific type of observer model. In contrast, our approach minimizes a predictability surrogate that allows modeling the observers with an arbitrary prediction model choice.
@@ 4.1. Derivation of a predictability aware cost function
where $\mathbb{KL}$ denotes the KL-Divergence. The right-hand side provides an upper bound on the free energy, and one can minimize this instead of the free energy. It resembles a standard control objective, and the terms allow for good conceptual understanding of the effect they have: A Performance Cost and Predictability Cost respectively, which penalizes agents for acting unpredictably. Using this newly found expression as a stage cost, we can craft the following cost function as a stage cost for a planning problem:
@@
We can emphasize now the intuition behind using the free energy as a way of incorporating predictability into optimal control. Eq. (4) is minimized precisely when $\mathcal{Q}=\mathcal{Q}^{*}=\mathcal{P}$. That is, the trajectory distribution executed is exactly the optimal cost trajectory distribution, and this matches the predicted distribution. Under this condition, the agent is behaving without surprising external observers and simultaneously obtaining optimal cost in its objective.
@@ 4.2. Integration with a Planner and Practicalities
$J(\tau_{0:K})=\sum_{k=0}^{K}J_{k}(\boldsymbol{x}_{k},\boldsymbol{u}_{k})+\gamma^{k}\lambda\mathbb{KL}(q(\boldsymbol{x}_{k})||p(\boldsymbol{x}_{k})).$ (5)
Remark 0.
It should be noted that our method is agnostic to the choice of the planner. However, in case MPPI is used as the planner, similar to Trevisan and Alonso-Mora (2024), our approach can be interpreted as leveraging the distribution of the prediction model as an ancillary controller to influence the MPPI sampling process.
@@ 5.1. Planner
For all experiments in this section, we use a sampling-based planner, namely Model Predictive Path Integral (MPPI) control Pezzato et al. (2023), based on the methodology presented in Williams et al. (2017b). MPPI places no restrictions on dynamics model or cost function and converges well toward optima with a moderate amount of samples Williams et al. (2017a). Given a nominal control sequence as an initial guess, MPPI applies Gaussian noise at each step to generate a set of $M$ control sequence samples. It then uses a state transition function $f(\cdot)$ to simulate their corresponding $M$ state trajectories. Each of the resulting state trajectories is evaluated based on the cost defined in (5), resulting in a total sample cost $J_{m}$. Once $J_{m}$, $\forall m\in[1,...,M]$ is computed, importance sampling weights, $w_{m}$, can be calculated as:
@@ 5.4.2. Robot-Robot Traffic Scenario
When agents fail to coordinate in road scenarios, they often experience deadlocks or, in the worst case, collisions. In Figure 3(a), an example of a deadlock is illustrated. Deadlocks are common in limited space environments such as intersections or narrow passages. Initially, the model may predict one agent will yield while the other advances. However, as deviations occur and both agents hesitate, their predictions begin to reinforce each other’s hesitation, creating the deadlock. The model may then be unable to introduce asymmetry to prioritize one of the agents in ambiguous situations, preventing the agents from breaking away from the deadlock. Results show that agents incorporating predictability into their models achieve better coordination, as indicated by less pronounced slowdowns resulting in higher traveled distance and the disappearance of deadlocks as seen in Table 2. When examining other metrics, the benefits of incorporating predictability are not as pronounced, especially for higher $\lambda$. This occurs because the prediction model is not explicitly conditioned to align with the road geometry (Figures 3(c),3(d)). Since the planner is required to track a reference path, deviations between the predictions and the reference path can push the agent to deviate from the path, requiring small adjustments more frequently for higher higher $\lambda$. Although this problem has marginal impact on the overall performance of the agent, the reduction in planning effort may be mitigated.
@@ Discussion
Balancing predictability and performance cost, determined by $\lambda$, is complex and context-dependent. Dynamically adjusting $\lambda$ as agents interact could improve performance, increasingly prioritizing predictability in safety-critical moments. Developing adaptive heuristics for this adjustment, as suggested by previous work Bastarache et al. (2023); Dragan and Srinivasa (2014), would be a valuable research direction. Alternatively, using lexicographic optimization could enhance generalizability by providing a structured trade-off that eliminates the need for tuning a magnitude dependent weighting parameter. However, This requires adpatation of the cost function computation to account for the lexicographic priorities, where predictability is prioritized subject to a performance constraint.